OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the following system. Can you please show me step by step? I'm really confused and I still haven't been able to find the answer. 3x – 4y = –7 x + y = 7

OpenStudy (zpupster):

find the denominator first: take the coefficients of x and y |3 -4| |1 1| cross multiply 3 x 1 and 1 x -4 = 3 -(-4) = 3+4 = 7 to get the numerator next substitute the x coefficients with the constants on right had side of equations |-7 -4| | 7 1| cross multiply like before -7 - (-28) = 21 then 21/7 = 3 this our actual x lets solve for y x+y = 7 3 + 4 = 7 and in the other equation 3(3) -4(-4) = -7 9-16 = -7

OpenStudy (anonymous):

@zpupster thank you for responding! That's what I got too, but I'm still confused because the only choices I have to choose from are: 21, 28, -14 or -35. ???

OpenStudy (zpupster):

21 is you DX

OpenStudy (zpupster):

we got carried away and solved, it is just asking for DX

OpenStudy (anonymous):

@zpupster you're my hero! Thanks so much :~)